# MPCM-Net: Multi-scale network integrates partial attention convolution with Mamba for ground-based cloud image segmentation

### v1

@@ Abstract
Ground-based cloud image segmentation is a critical research domain for photovoltaic (PV) power forecasting. Current deep learning approaches primarily focus on encoder-decoder architectural refinements. However, existing methodologies exhibit several limitations: (1) they rely on dilated convolutions for multi-scale context extraction, lacking the partial feature effectiveness and interoperability of inter-channel; (2) attention-based feature enhancement implementations neglect accuracy-throughput balance; and (3) the decoder modifications fail to establish global interdependencies among hierarchical local features, limiting inference efficiency. To address these challenges, we propose MPCM-Net, a Multi-scale network that integrates Partial attention Convolutions with Mamba architectures to enhance segmentation accuracy and computational efficiency. Specifically, the encoder incorporates multi-scale partial attention convolution (MPAC), which comprises: (1) a multi-scale partial convolution block (MPC) with partial channel module (ParCM) and partial spatial module (ParSM) that enables global spatial interaction across multi-scale cloud formations, and (2) a multi-scale partial attention block (MPA) combining partial attention module (ParAM) and ParSM to extract discriminative features with reduced computational complexity. On the decoder side, a multi-scale Mamba block (M2B) is employed to mitigate contextual loss through a spatial-semantic hybrid domain (SSHD) that maintains linear complexity while enabling deep feature aggregation across spatial and scale dimensions. As a key contribution to the community, we also introduce and release a dataset incorporating Complex-Scale variations, Radiative properties, and Color attributes (CSRC), which is a clear-label, fine-grained segmentation benchmark designed to overcome the critical limitations of existing public datasets. Extensive experiments on CSRC demonstrate the superior performance of MPCM-Net over state-of-the-art methods, achieving an optimal balance between segmentation accuracy and inference speed. The dataset and source code will be available at https://github.com/she1110/CSRC.
@@ I Introduction
Clouds constitute critical atmospheric components governing solar radiation transmission and energy balance, directly determining surface insolation intensity with profound implications for agricultural irrigation, aviation safety, and renewable energy forecasting [bg]. Contemporary cloud monitoring technologies are broadly classified into satellite remote sensing [satellite] and ground-based observations [PV], with the latter offering superior spatiotemporal resolution (updating at second-scale intervals) and minimized atmospheric attenuation at low observation altitudes. This enables the precise capture of cloud morphology, textural features, and kinematic trajectories. Particularly in PV power forecasting, precisely segmented ground-based cloud imagery facilitates real-time identification of cloud formations, providing minute-resolution irradiance fluctuation alerts for solar farms. Consequently, it significantly improves power generation forecast accuracy while delivering unique value for grid dispatch operations and clean energy integration.
@@
Advancing the state-of-the-art of accurate ground-based cloud image segmentation requires overcoming two core technical challenges: a) developing models with the capacity to character multi-scale morphological variations within individual cloud formations across sequential frames, and b) implementing computationally efficient algorithms that precisely capture intricate boundary details while satisfying real-time inference constraints for operational deployment.
@@
Existing methods have sought to address these challenges through two primary avenues: approaches based on image processing with handcrafted features, and the more recently emerging deep learning-based methods. The former category, which includes threshold-based [threshold] and spatial feature-based approaches [spatial], often suffers from being computationally intensive and requiring extensive manual annotation, resulting in suboptimal segmentation of cloud imagery details. Methods that rely on engineered features are often hindered in their ability to capture the multi-scale characteristics of cloud formations, thereby restricting model accuracy and generalization ability [ML].
@@
In contrast, deep learning (DL) techniques have garnered significant attention due to their powerful feature extraction capabilities. Numerous studies have validated the effectiveness of classical semantic segmentation algorithms, such as U-Net [U-Net], DeepLab [DeepLabv3], and SegNet [SegNet], in ground-based cloud image segmentation tasks [CloudSegNet, SegCloud]. To extract features from multi-scale cloud formations, some researchers have proposed incorporating multi-scale dilated convolutions within the encoder to expand the model’s receptive fields, thereby enhancing the network’s multi-scale feature extraction capacity [CloudU-Net, MA-SegCloud]. Furthermore, some researchers have introduced attention mechanisms into the encoder; by dynamically re-weighting features to focus on the most salient information, these mechanisms are adept at capturing long-range contextual dependencies, thereby improving multi-scale feature representation [Att1, Att2]. To mitigate contextual information loss and alleviate inaccuracies in segmenting boundary details, several scholars have proposed enhancements to upsampling operations [CloudU-Netv2, CloudDeepLabV3+]. Because standard upsampling methods can introduce artifacts or fail to recover fine-grained details lost during encoding, these advanced, often learnable operators aim to reconstruct high-resolution feature maps with greater fidelity. Concurrently, auxiliary modules have been developed to integrate multi-scale features within the decoder, compensating for local information loss during upsampling [TransCloudSeg, BSANet]. Self-attention mechanisms, widely employed for capturing long-range dependencies, have also proven effective in addressing contextual loss [ViT, SViT]. Some researchers have integrated various attention mechanisms into the decoder to strengthen global feature representation [CloudDeepLabV3+, MA-SegCloud].
@@
Despite these encoder and decoder refinements improving ground-based cloud image segmentation performance, several limitations persist. Firstly, the multi-scale dilated convolution methods employed in encoders often overlook the efficacy and intercommunity of partial spatial and channel features, impairing global information interaction. Moreover, the implemented attention mechanisms frequently disregard the computational overhead and latency implications for the network model. Furthermore, decoder enhancements commonly neglect the correlation between global features and local information across different hierarchical levels, as well as the critical trade-off between segmentation accuracy and inference efficiency.
@@
Accurate segmentation of ground-based cloud images represents a critical task in PV power forecasting. However, most publicly available datasets for ground-based cloud image segmentation currently only support binary segmentation between clouds and sky background, which hinders the accurate prediction of PV power generation. This limitation primarily originates from the fact that radiation sources and cloud layers of varying colors severely restrict the utility of these segmentation maps for accurate, real-world PV power forecasting, which is highly dependent on cloud optical properties. Consequently, the development of fine-grained cloud image segmentation datasets and techniques that incorporate radiation sources and color categorization is particularly imperative.
@@
To address these distinct challenges, this paper proposes the Multi-scale network integrates partial attention convolution with Mamba (MPCM-Net), a novel architecture that integrates multi-scale partial attention convolution in the encoder and multi-scale Mamba in the decoder. Specifically, at the encoder stage, we introduce a multi-scale partial attention convolution (MPAC) to adaptive extract multi-scale contextual information from ground-based cloud imagery while enhancing feature interaction. The MPAC comprises two key components: 1) a multi partial convolution block (MPC) incorporating partial channel module (ParCM) and partial spatial module (ParSM) is designed to facilitate interaction between partial and global information across diverse cloud formations; and 2) a multi partial attention block (MPA) combining partial attention module (ParAM) and ParSM, which employs a novel attention mechanism to expand the network’s receptive field and improve multi-scale feature perception. For the decoder part, a multi-scale Mamba module (M2B) is implemented to mitigate boundary detail loss during upsampling. M2B introduces a spatial-semantic hybrid domain (SSHD) that enhances feature representation across spatial and cross-scale dimensions while maintaining linear computational complexity, thereby achieving an effective balance between segmentation accuracy and inference speed. Finally, we develop and release a new dataset that incorporates fine-grained segmentation based on radiative sources and color attributes. The main contributions of this article are summarized as follows.
@@
    A novel multi-scale network for ground-based cloud image segmentation, MPCM-Net, is proposed. It explicitly integrates multi-scale partial attention convolution with Mamba architecture to enhance multi-scale feature extraction, thereby strengthening boundary feature representation while significantly improving segmentation accuracy and computational efficiency.
@@
    A multi-scale partial attention convolutional (MPAC) is introduced in the encoder. It features a multi partial convolution block that leverages partial channel and spatial information to facilitate global information exchange across diverse cloud formations and a multi partial attention block to optimize multi-scale feature perception while maintaining computational efficiency.
@@
    The introduction and public release of the CSRC dataset. As a major contribution to the community, the CSRC dataset is a newly introduced complex-scale, clear-label, and fine-grained segmentation benchmark that addresses critical limitations of prior datasets. It establishes a more realistic and challenging foundation for advancing ground-based cloud image segmentation research. Extensive experiments on CSRC demonstrate that MPCM-Net outperforms state-of-the-art methods in both prediction accuracy and computational efficiency.
@@ II Related Works
Accurate segmentation of ground-based cloud imagery holds critical significance for climate change studies, environmental monitoring, and weather forecasting, attracting considerable research attention in recent years. Current methodologies primarily bifurcate into traditional-based methods and deep learning-based methods. Traditional algorithms encompass threshold-based approaches, spatial feature-based approaches, and machine learning-based approaches. However, threshold-based methods exhibit insufficient precision for complex scenarios, while spatial feature-based and machine learning-based approaches suffer from limited accuracy and generalizability due to dependency on hand-crafted feature engineering.
With rapid advances in deep learning for computer vision, researchers have achieved remarkable performance in ground-based cloud image segmentation by employing a DL architecture. Currently, most DL-based methods use an encoder-decoder architecture, where both components are continually enhanced to address domain-specific technical challenges. In addition, many scientific research units and research institutions use advanced all-sky imagers to obtain high-resolution cloud images and construct datasets.
@@ II-A Encoder for Ground-based Cloud Image Segmentation
The core function of an encoder involves extracting features and compressing information from input data, transforming raw inputs into semantically rich representations. U-Net, introduced by Ronneberger et al. for biomedical image segmentation, has been widely adopted in semantic segmentation research [U-Net]. Its encoder typically employs conventional CNNs like VGG-16 or ResNet-50 for feature extraction. Gonzales et al. proposed pre-trained deep U-Net networks for cloud layer segmentation, demonstrating superior performance [Gonzales]. Lu et al. proposed a symmetric network based on SegNet [SegNet] for cloud segmentation accuracy [Lu]. Chen et al. further advanced encoder design through DeepLabv3+, incorporating depthwise separable convolutions to enhance segmentation efficiency and precision [DeepLabv3]. Strengthening the encoder’s capacity for multi-scale contextual feature extraction is critical for improving ground-based cloud image segmentation accuracy. Shi et al. integrated dilated convolutions into the U-Net’s encoder to expand receptive fields and capture multi-scale features [CloudU-Net]. Zhang et al. incorporated a Multi-branch Asymmetric Convolution Module (MACM) in the encoder, enhancing the model’s capability to extract multi-scale contextual information through adaptive channel weighting and receptive field adjustment [MA-SegCloud]. Li et al. employed depthwise (DW) convolution and a Heterogeneous Receptive Field-enhanced ASPP module (HRFS-ASPP) in the encoder to extract multi-scale features while enhancing inter-feature relevance [CloudDeepLabV3+]. Shi et al. incorporated an adaptive kernel selection mechanism in the encoder to extract multi-scale features from cloud imagery, complemented by a parallel dilated convolution module that refines multi-scale feature representation [CloudFU-Net]. Li et al. proposed a Dynamic Multi-scale Convolution (DMSC) module in the encoder that adaptively expands the receptive field to enhance the network’s multi-scale feature perception capability [DDUNet]. Hu et al. enhanced cloud region recognition capabilities by incorporating embedding layers within the encoder [FA-CloudSeg]. Feng et al. proposed the Hybrid Strip Pooling Module (HSPM) in the encoder to enhance its capability for extracting fine-grained local features; this module employs a dual-branch strip pooling operation to minimize local information loss [MFAFNet]. With attention mechanisms becoming increasingly prevalent across computer vision tasks, researchers have adapted these approaches for ground-based cloud image segmentation [SViT, ViT]. Shi et al. introduced a Fine-grained Feature Fusion Module (Fg-FFM) in the encoder with self-attention to extract multi-scale features while integrating contextual information [CloudSwinNet]. Liu et al. employ a hybrid CNN-Transformer design in the encoder to extract features across varying scales and properties [TransCloudSeg]. Guo et al. introduced a channel attention mechanism in the U-Net encoder to improve feature representation ability [Guo].
However, these methods merely fuse multi-scale features through simple channel concatenation in their encoder-based dilated pyramidal structures, overlooking the efficacy of partial spatial-channel feature interaction and compromising global information integration. Moreover, most encoders employing attention mechanisms neglect computational load and inference latency impacts, thus hindering optimal balance between segmentation accuracy and inference speed in ground-based cloud image segmentation.
@@ II-B Decoder for Ground-based Cloud Image Segmentation
The primary function of the decoder is to progressively generate task-aligned outputs from the semantically enriched feature vectors of the encoder. Upsampling operations within decoders, however, incur partial contextual information loss, resulting in incomplete boundary segmentation for ground-based cloud imagery. Shi et al. used a bilinear upsampling operation instead of the standard upsampling operation in the original U-Net decoder to alleviate the problem of context loss [CloudU-Netv2]. Multi-scale feature fusion and auxiliary module integration significantly enhance decoder feature representation. Shi et al. introduced content-aware reassembly of features (CARAFE) to mitigate information loss during upsampling [CloudFU-Net]. Li et al. introduced a dynamic weights and bias generator (DWBG) module that incorporates features from both the encoder and decoder to improve model generalization [DDUNet]. Li et al. proposed a Bilateral Segregation and Aggregation Module (BSAM) that integrates coarse segmentation results derived from ground truth labels with features during the decoder upsampling stage, thereby improving segmentation accuracy [BSANet]. Feng et al. introduced the Three-Branch Adaptive Feature Fusion Module (TB-AFFM) to effectively fuse multi-scale features from the encoder during upsampling, mitigating contextual loss of information [MFAFNet]. Employing attention mechanisms to capture long-range dependencies in cloud clusters effectively preserves boundary integrity in ground-based cloud images. Zhang and Feng et al. integrated the convolutional block attention module (CBAM) in the decoder to strengthen feature restoration during upsampling [MA-SegCloud, MFAFNet]. Li et al. utilized a heterogeneous fusion module in the decoder that leverages self-attention mechanisms to integrate features from different encoder stages, thereby enhancing feature representational capacity [TransCloudSeg]. Li and Hu et al. proposed the feature alignment module in the decoder that leverages self-attention mechanisms to integrate multi-scale contextual information [CloudDeepLabV3+, FA-CloudSeg]. Shi et al. employed a Global Graph Reasoning (GloRe) module that establishes an interaction space to achieve mutual reinforcement between global and local features, thereby enhancing the model’s capacity for learning long-range dependencies [CloudSwinNet]. Wei et al. propose a Token Dual Attention module (TDA) bridging the encoder and decoder, which enhances feature discriminability through pooled convolutions and coordinate attention mechanisms, thereby mitigating contextual information loss in the decoder [DASUNet].
Unfortunately, these approaches fail to capture cross-hierarchical global feature correlations during upsampling and substantially increase computational complexity when employing self-attention-like mechanisms, thereby compromising inference efficiency in ground-based cloud image segmentation.
Figure 1: (a) illustrates multi-scale cloud. The red, yellow, and green blocks represent large, medium, and small-scale clouds clusters, respectively. In adjacent frames, the same cloud presents different scales. (b) illustrates the incomplete extraction of the boundary of clouds. Note that the results are from U-Net.

### v2

# MPCM-Net: A multi-scale network that integrates partial attention convolution with Mamba for ground-based cloud image segmentation

@@ Abstract
Ground-based cloud image segmentation is a critical research domain for photovoltaic (PV) power forecasting. Current deep learning (DL) approaches primarily focus on encoder-decoder architectural refinements. However, existing methodologies exhibit several limitations: (1) they rely on dilated convolutions for multi-scale context extraction, yet fail to leverage inter-channel interoperability and partial feature efficacy; (2) implementations of attention-based feature enhancement frequently compromise the equilibrium between accuracy and throughput; and (3) the decoder modifications often fail to re-establish global interdependencies among hierarchical local features, thereby constraining inference efficiency. To mitigate these challenges, we propose MPCM-Net, a Multi-scale network that integrates Partial attention Convolutions with Mamba architectures to enhance segmentation accuracy. Specifically, the encoder incorporates a multi-scale partial attention convolution (MPAC), which comprises: (1) a multi-scale partial convolution block (MPC) with partial channel module (ParCM) and partial spatial module (ParSM) that facilitating global spatial interaction across multi-scale cloud formations, and (2) a multi-scale partial attention block (MPA) combining partial attention module (ParAM) and ParSM to extract discriminative features with reduced computational complexity. On the decoder side, a multi-scale Mamba block (M2B) is employed to mitigate contextual loss through a spatial-semantic hybrid domain (SSHD) that maintains linear complexity while enabling deep feature aggregation across spatial and scale dimensions. Furthermore, we introduce and release a dataset incorporating Complex-Scale variations, Radiative properties, and Color attributes (CSRC), which is a clear-label, fine-grained segmentation benchmark designed to overcome the critical limitations of existing public datasets. Extensive empirical analysis on CSRC demonstrates the superior performance of MPCM-Net over state-of-the-art methods, achieving an optimal balance between segmentation accuracy and inference speed. The dataset and source code will be available at https://github.com/she1110/CSRC.
@@ I Introduction
Clouds constitute critical atmospheric components governing solar radiation transmission and energy balance, directly determining surface insolation intensity with profound implications for agricultural irrigation, aviation safety, and renewable energy forecasting [19]. Contemporary cloud monitoring technologies are broadly categorized into satellite remote sensing [47] and ground-based observations [43], with the latter offering superior spatiotemporal resolution (updating at second-scale intervals) and minimized atmospheric attenuation at low observation altitudes. This enables the precise capture of cloud morphology, textural features, and kinematic trajectories. Particularly in PV power forecasting, precisely segmented ground-based cloud imagery facilitates real-time identification of cloud formations, providing minute-resolution irradiance fluctuation alerts for solar farms. Consequently, it significantly improves power generation forecast accuracy while delivering unique value for grid dispatch operations and clean energy integration.
@@
Advancing the state-of-the-art (SOTA) of accurate ground-based cloud image segmentation requires overcoming two core technical challenges: a) developing models with the capacity to characterize multi-scale morphological variations within individual cloud formations across sequential frames, and b) implementing computationally efficient algorithms that precisely capture intricate boundary details while satisfying real-time inference constraints for operational deployment.
@@
Existing methods have sought to address these challenges through two primary avenues: approaches based on image processing with handcrafted features, and the more recently emerging DL-based methods. The former category, which includes threshold-based [37] and spatial feature-based approaches [50], typically entails high computational costs and requires extensive manual annotation, resulting in suboptimal segmentation of cloud imagery details. Methods that rely on engineered features are often hindered in their ability to capture the multi-scale characteristics of cloud formations, thereby restricting model accuracy and generalization ability [5].
@@
In contrast, DL methodologies have gained prominence due to their superior automated feature extraction capabilities. Classical semantic segmentation architectures, such as U-Net [36], DeepLab[4], and SegNet[1], have demonstrated significant efficacy in ground-based cloud segmentation tasks[7, 46]. To specifically address the multi-scale nature of cloud formations, researchers have integrated multi-scale dilated convolutions within encoders, effectively expanding receptive fields without loss of resolution[40, 48]. Concurrently, attention mechanisms have been introduced to dynamically re-weight features; by prioritizing salient information, these mechanisms facilitate the capture of long-range contextual dependencies, thereby refining multi-scale feature representation[2, 17]. Ideally, the decoder should recover spatial details lost during the encoding process. However, standard upsampling operators often introduce artifacts or fail to reconstruct fine-grained boundary details. To mitigate this contextual information loss, significant research has focused on advanced, learnable upsampling operators[41, 27]. Simultaneously, auxiliary modules and Self-Attention (SA) mechanisms have been integrated into decoders to aggregate multi-scale features and reinforce global representations[10, 32].
@@
Despite these encoder and decoder refinements improving segmentation performance, several limitations persist. Firstly, the multi-scale dilated convolution methods employed in encoders often overlook the efficacy and interconnectivity of partial spatial and channel features, impairing global information interaction. Moreover, the implemented attention mechanisms frequently disregard the computational overhead and latency implications for the network model. Furthermore, decoder enhancements commonly neglect the correlation between global features and local information across different hierarchical levels, as well as the critical trade-off between segmentation accuracy and inference efficiency.
@@
Accurate segmentation of ground-based cloud images represents a critical task in PV power forecasting. However, most publicly available datasets currently only support binary segmentation between clouds and sky background, which hinders the accurate prediction of PV power generation. This limitation primarily originates from the fact that radiation sources and cloud layers of varying colors severely restrict the utility of these segmentation maps for accurate, real-world PV power forecasting, which is highly dependent on cloud optical properties. Consequently, the development of fine-grained cloud image segmentation datasets and techniques that incorporate radiation sources and color categorization is particularly imperative.
@@
To address these distinct challenges, this paper proposes the Multi-scale network that integrates partial attention convolution with Mamba (MPCM-Net), a novel architecture that integrates multi-scale partial attention convolution in the encoder and multi-scale Mamba in the decoder. Specifically, at the encoder stage, we introduce a multi-scale partial attention convolution (MPAC) to adaptively extract multi-scale contextual information while enhancing feature interaction. The MPAC comprises two key components: 1) a multi partial convolution block (MPC) incorporating partial channel module (ParCM) and partial spatial module (ParSM) is designed to facilitate interaction between partial and global information across diverse cloud formations; and 2) a multi partial attention block (MPA) combining partial attention module (ParAM) and ParSM, which employs a novel attention mechanism to expand the network’s receptive field and improve multi-scale feature perception. For the decoder part, a multi-scale Mamba module (M2B) is implemented to mitigate boundary detail loss during upsampling. M2B introduces a spatial-semantic hybrid domain (SSHD) that enhances feature representation across spatial and cross-scale dimensions, thereby achieving an effective balance between segmentation accuracy and inference speed. Finally, we develop and release a new dataset that incorporates fine-grained segmentation based on radiative sources and color attributes. The main contributions of this article are summarized as follows.
@@
    A novel multi-scale network for ground-based cloud image segmentation, MPCM-Net, is proposed. It explicitly integrates multi-scale partial attention convolution with Mamba architecture to enhance multi-scale feature extraction, thereby strengthening boundary feature representation while significantly improving segmentation accuracy.
@@
    A multi-scale partial attention convolutional is introduced in the encoder. It features a multi partial convolution block that leverages partial channel and spatial information to facilitate global information exchange across diverse cloud formations and a multi partial attention block to optimize multi-scale feature perception while maintaining computational efficiency.
@@
    The introduction and public release of the CSRC dataset. As a major contribution to the community, the CSRC dataset is a newly introduced complex-scale, clear-label, and fine-grained segmentation benchmark that addresses critical limitations of prior datasets. It establishes a more realistic and challenging foundation for advancing ground-based cloud image segmentation research. Extensive experiments on CSRC demonstrate that MPCM-Net outperforms SOTA methods in both prediction accuracy and computational efficiency.
@@ II Related Works
Accurate segmentation of ground-based cloud imagery holds critical significance for climate change studies, environmental monitoring, and weather forecasting, garnering significant scholarly interest in recent years. Methodologies in this domain are generally categorized into traditional algorithms and DL-based methods. Traditional algorithms, encompassing threshold-based, spatial feature-based, and machine learning-based approaches, are frequently constrained by limited accuracy and generalizability due to their dependency on hand-crafted feature engineering. Conversely, with rapid advances in computer vision, DL-based methods have achieved remarkable performance by employing encoder-decoder architectures that automatically learn hierarchical representations.
@@ II-A Encoder for Ground-based Cloud Image Segmentation
The encoder fundamentally serves to extract features and compress information, transforming raw inputs into semantically rich representations. While foundational architectures like U-Net[36] and DeepLabv3+[4] have been widely adopted, recent research focuses on enhancing the encoder’s capacity to handle the extreme scale variations and complex boundaries of cloud formations [14, 33]. Strengthening the encoder’s ability to capture multi-scale features is paramount. Early approaches integrated dilated convolutions into U-Net-like structures to expand receptive fields without losing spatial resolution [40, 29]. To accommodate the irregular geometric morphology of clouds, Zhang et al. introduced multi-branch asymmetric convolution modules [48]. These mechanisms utilize non-square convolution kernels to better capture directional features and employ adaptive channel weighting to refine contextual information. Similarly, Li et al. have proposed heterogeneous receptive field enhancement strategies, which combine depthwise separable convolutions with varied dilation rates to extract dense multi-scale features efficiently [27]. More recently, dynamic multi-scale convolution mechanisms constitute a significant advancement; in contrast to static filters, these modules adaptively expand the receptive field based on the input image properties, significantly improving the network’s perception of variable cloud structures [39]. Standard convolutional operations often struggle to model long-range dependencies in large-scale cloud systems. To mitigate this, strip pooling operations have been adopted to capture global context along narrow axes, which is particularly effective for segmenting streak-like cloud patterns [13]. Additionally, Hu et al. explored the incorporation of embedding layers to enhance regional recognition capabilities [22]. With the prevalence of attention mechanisms, hybrid architectures integrating CNNs and Transformers have emerged. Shi et al. have designed fine-grained feature fusion modules to integrate local details with global context [38]. Hybrid CNN-Transformer encoders allow for the simultaneous extraction of features across varying scales and properties [30]. Furthermore, Guo et al. proposed the channel attention mechanisms that are frequently embedded within encoder blocks to recalibrate feature responses [18].
However, many of these methods rely on simple concatenation for fusion, potentially overlooking spatial-channel interactions and incurring high computational costs that hinder real-time deployment.
@@ II-B Decoder for Ground-based Cloud Image Segmentation
The primary function of the decoder is to progressively generate task-aligned outputs from the enriched feature vectors. A critical challenge here is recovering the spatial details lost during downsampling. Standard upsampling operations often result in blurred boundaries. To address this, Shi et al. have introduced bilinear upsampling and, more notably, content-aware feature reassembly operators. These advanced operators dynamically predict reassembly kernels based on local content, significantly reducing information loss and sharpening boundary delineation [41, 39]. Moreover, to improve generalization across diverse atmospheric conditions, Li et al. proposed dynamic weight generation mechanisms in the final classification layers, allowing the model to adaptively adjust its decision boundaries based on aggregated encoder-decoder features [29]. Effective fusion of high-level semantics and low-level details is essential for precise segmentation. Bilateral segregation and aggregation strategies have been proposed to leverage coarse segmentation priors to guide the refinement of fine-grained features [28]. Similarly, Feng et al. constructed multi-branch adaptive fusion units to intelligently merge features from different network depths, balancing local and global information [13]. To further strengthen boundary integrity, attention mechanisms are extensively applied in the decoder. Convolutional block attention modules are frequently integrated to enhance feature restoration during upsampling [48, 13]. For modeling complex, non-local relationships between disjoint cloud regions, Shi and Liu et al. have employed global graph reasoning frameworks. These methods project features into an interaction space to infer long-range dependencies before mapping them back to the pixel space [30, 38]. Additionally, token-based dual attention mechanisms have been utilized in skip connections to bridge the encoder and decoder, enhancing feature discriminability. Despite these innovations, the computational complexity of such attention mechanisms remains a concern for inference efficiency [48, 44].
Notwithstanding these innovations, these approaches fail to capture cross-hierarchical global feature correlations during upsampling, and the computational complexity of such attention mechanisms remains a concern for inference efficiency.